# Fast Estimation of the Composite Link Model for Multidimensional Grouped Counts

### v1

@@ 3.2 Mortality grouped by age over years and weeks
For this application, we present an actual dataset for Spanish males in which mortality is analyzed by age, year, and week. Specifically, data on deaths are available by age group (with intervals of 5 years: 0-4, 5-9, $\ldots$, 85-89, 90+) and across the years 2000 to 2019, as well as by week (1-52). These death counts were sourced from the Instituto Nacional de Estadística (2024) and provided by the Short-Term Mortality Fluctuations data series (2024).
The dataset for the exposures was obtained from the Human Mortality Database (2024), where data is available by single years of age, ranging from 0 to 104 years, and for each year. To ensure the temporal consistency of the dataset for analysis, death counts were adjusted to align with a 52-week year, and the original annual exposure data were linearly interpolated to achieve a weekly resolution.
@@
With our approach, these matrices do not need to be explicitly constructed. This eliminates the associated storage and computational challenges, allowing the model to be fitted efficiently and without memory constraints. Notably, the entire R workspace, which includes all the data and resulting outcomes, occupies only 147 MB.
Despite the model’s complexity, as indicated by the large number of penalized parameters ($\bm{\alpha}$, equal to $c_{1}c_{2}c_{3}=840$), the adoption of the GLAM arithmetic enables efficient estimation within a reasonable computational time: approximately 29 seconds under the model settings and the previously described hardware configuration. Notably, in this three-dimensional case, a substantial portion of this time (45%) is dedicated to computing the model uncertainty, a step where the full GLAM algorithm cannot be utilized (see Section 2.3). This percentage is notably higher than in the two-dimensional case, reflecting the increased computational demands of the higher-dimensional model.
@@ 4 Conclusions
In fields like demography, epidemiology, or economics, data aggregation over different dimensions such as age or time is often unavoidable due to privacy concerns or data collection constraints. Such aggregation can obscure latent patterns that are crucial for understanding underlying processes and informing policy decisions. The penalized version of the Composite Link Model model enables the disaggregation of grouped data to capture fine-grained trends in mortality and other datasets, offering practical utility for applications requiring detailed insights. However, model estimation becomes computationally demanding, even impossible, since memory and processing demands can become prohibitive when the dimension of the array and/or the desired refinement of the latent distribution to be estimated increases.
To address this issue, we proposed a modified version of the original algorithm introduced by Thompson and Baker (1981), reformulating the PCLM estimation process. This reformulation is based on defining a working latent response rather than a working regression matrix. This is similar to the E-step in the EM algorithm, but is done purely for computational convenience. Unlike the EM algorithm, which assumes a distribution for the latent data, our approach assumes a Poisson distribution for the observed aggregated counts. This key distinction ensures that the method focuses directly on the observed data without alternating between observed and latent distributions. By redefining the algorithm in terms of the working response, the entire process is reformulated as a Generalized Linear Array Model (GLAM, Currie et al., 2006), significantly reducing computational complexity and enabling the estimation of models that were previously computationally infeasible.
The presented applications demonstrate the versatility and computational efficiency of the proposed methodology using two real-world mortality examples. First, Swedish female mortality grouped by age and year is analyzed, showcasing the method’s capability to disaggregate aggregated counts into finer resolutions while preserving accuracy and computational feasibility. Second, a more complex dataset of Spanish male mortality, classified by age group, year, and week, highlights the methodology’s extension to three-dimensional settings. These applications underscore the practicality of the approach for handling large-scale, multidimensional grouped observations, demonstrating its potential for generating fine-grained insights from aggregated observations.
@@ Appendix A Computational Implementation
As mentioned in the main body of the document, Currie et al. (2006) developed an arithmetic for arrays that enables high-speed computation in the scoring algorithm of generalized linear models. This arithmetic is based on a sequence of nested matrix operations, such that by reorganizing the computations, it becomes possible to perform operations using arrays of the same size as the data. This approach avoids the need to vectorize the data on the array or flatten the tensor products to a regression basis, resulting in significant computational gains as the array’s dimensionality increases. Here, We outline the fundamental operations required to solve (11) and provide their implementation in R code. The fully reproducible program, which estimates the models for both datasets and reproduces the plots presented in this paper, is available at osf.io/uwejt/?view_only=2ca1fdb7568342bbb9a3c51fd33c718c (temporarily anonymized to maintain confidentiality during the review process).

### v2

@@ 4.2 Mortality grouped by age over years and weeks
For this application, we present an actual dataset for Spanish males in which mortality is analyzed by age, year, and week. Specifically, data on deaths are available by age group (with intervals of 5 years: 0-4, 5-9, $\ldots$, 85-89, 90+) and across the years 2000 to 2019, as well as by week (1-52). These death counts were sourced from the INEweekData and provided by the STMF2024.
The dataset for the exposures was obtained from the HMD, where data are available by single years of age, ranging from 0 to 104 years, and for each year. To ensure the temporal consistency of the dataset for analysis, death counts were adjusted to align with a 52-week year, and the original annual exposure data were linearly interpolated to achieve a weekly resolution.
@@
With our approach, these matrices do not need to be explicitly constructed. This eliminates the associated storage and computational challenges, allowing the model to be fitted efficiently and without memory constraints. Notably, the combined size of all relevant R objects, including the data and resulting outcomes, amounts to only 319 MB.
Despite the model’s complexity, as indicated by the large number of penalized parameters ($\bm{\alpha}$, equal to $c_{1}c_{2}c_{3}=840$), the adoption of the GLAM arithmetic enables efficient estimation within a reasonable computational time: approximately 25 seconds under the model settings and the previously described hardware configuration. Notably, in this three-dimensional case, a substantial portion of this time (45%) is dedicated to computing the model uncertainty, a step where the full GLAM algorithm cannot be utilized (see Section 2.3). This percentage is notably higher than in the two-dimensional case, reflecting the increased computational demands of the higher-dimensional model.
@@ 5 Conclusions
In fields like demography, epidemiology, or economics, data aggregation over different dimensions such as age or time is often unavoidable due to privacy concerns or data collection constraints. Such aggregation can obscure latent patterns that are crucial for understanding underlying processes and informing policy decisions. The penalized version of the Composite Link Model model enables the disaggregation of grouped data to capture fine-grained trends in mortality and other datasets, offering practical utility for applications requiring detailed insights. However, model estimation becomes computationally demanding and sometimes even impossible, since memory and processing demands can become prohibitive when the dimension of the array and/or the desired refinement of the latent distribution to be estimated increases.
To address this issue, we proposed a modified version of the original algorithm introduced by ThompsonBakerCLM1981, reformulating the PCLM estimation process. This reformulation is based on defining a working latent response rather than a working regression matrix. This is similar to the E-step in the EM algorithm, but is done purely for computational convenience. Unlike the EM algorithm, which assumes a distribution for the latent data, our approach assumes a Poisson distribution for the observed aggregated counts. This key distinction ensures that the method focuses directly on the observed data without alternating between observed and latent distributions. By redefining the algorithm in terms of the working response, the entire process is reformulated as a Generalized Linear Array Model (GLAM, CurrieDurbanEilersGLAM2006), significantly reducing computational complexity and enabling the estimation of models that were previously computationally infeasible.
The simulation studies and real-world applications presented in this paper demonstrate the versatility and computational efficiency of the proposed methodology in the context of grouped count data. The first simulation study reaffirmed the high accuracy of the PCLM in recovering the latent distribution from aggregated data. The second study addressed the central contribution of this work, evaluating the performance of the new GLAM-based estimation algorithm across a variety of data structures, including two-, three-, and four-dimensional settings. The results clearly demonstrate that the proposed approach achieves substantial computational gains, by orders of magnitude, compared to the original PCLM algorithm; in fact, for large datasets, which are common in empirical applications, the original algorithm often fails to produce estimates altogether, making the proposed method not only more efficient but in many cases the only viable solution.
These findings are further supported by real-world applications. In particular, Swedish female mortality grouped by age and year is analyzed, illustrating the method’s capability to disaggregate aggregated counts into finer resolutions while preserving accuracy and computational feasibility. A more complex dataset of Spanish male mortality, classified by age group, year, and week, highlights the methodology’s extension to three-dimensional settings. These applications underscore the practicality of the approach for handling large-scale, multidimensional grouped observations, demonstrating its potential for generating fine-grained insights from aggregated observations.
@@ Appendix A Computational Implementation
As mentioned in the main body of the document, CurrieDurbanEilersGLAM2006 developed an arithmetic for arrays that enables high-speed computation in the scoring algorithm of generalized linear models. This arithmetic is based on a sequence of nested matrix operations, such that by reorganizing the computations, it becomes possible to perform operations using arrays of the same size as the data. This approach avoids the need to vectorize the data on the array or flatten the tensor products to a regression basis, resulting in significant computational gains as the array’s dimensionality increases. Here, We outline the fundamental operations required to solve (11) and provide their implementation in R code. The fully reproducible program, which estimates the models for both datasets and reproduces the plots presented in this paper, is available at osf.io/uwejt/?view_only=2ca1fdb7568342bbb9a3c51fd33c718c.